# Behind the Scenes: Mechanistic Interpretability of LoRA-adapted Whisper for Speech Emotion Recognition

### v2

@@ Abstract
Large pre-trained speech models such as Whisper offer strong generalization but pose significant challenges for resource-efficient adaptation. Low-Rank Adaptation (LoRA) has become a popular parameter-efficient fine-tuning method, yet its underlying mechanisms in speech tasks remain poorly understood. In this work, we conduct the first systematic mechanistic interpretability study of LoRA within the Whisper encoder for speech emotion recognition (SER). Using a suite of analytical tools, including layer contribution probing, logit-lens inspection, and representational similarity via singular value decomposition (SVD) and centered kernel alignment (CKA), we reveal two key mechanisms: a delayed specialization process that preserves general features in early layers before consolidating task-specific information, and a forward alignment, backward differentiation dynamic between LoRA’s matrices. Our findings clarify how LoRA reshapes encoder hierarchies, providing both empirical insights and a deeper mechanistic understanding for designing efficient and interpretable adaptation strategies in large speech models.
@@ 1 Introduction
Large-scale pre-trained models have fundamentally transformed speech and language processing. Models such as Whisper [1], Wav2Vec2 [2], and HuBERT [3] provide general-purpose representations that are widely adapted to diverse tasks, from automatic speech recognition (ASR) to SER. However, the heavy computational and storage demands of full fine-tuning hinder their deployment in resource-constrained settings. This has motivated the development of Parameter-Efficient Fine-Tuning (PEFT) methods [4, 5], which adapt these large encoders with minimal additional parameters.
Among PEFT approaches, LoRA [6] has demonstrated exceptional effectiveness across modalities including text, vision, and speech [7, 8]. By injecting low-rank trainable matrices into a frozen backbone, LoRA significantly reduces training costs while preserving the model’s powerful representational capacity. Despite its practical success, prior studies primarily emphasize performance gains, leaving a critical question unanswered: why does LoRA work so effectively?
SER serves as a compelling probe for this investigation. Unlike ASR, which primarily relies on lexical content, SER necessitates capturing subtle prosodic and paralinguistic cues that reflect high-level affective states [9]. This makes SER an ideal testbed for analyzing how LoRA reshapes a model’s internal representations and decision dynamics. Recent work has demonstrated the successful adaptation of both Whisper [10] and Wav2Vec2 [11] for SER, further underscoring the value of exploring parameter-efficient strategies in this domain.
In this work, we present the first systematic mechanistic analysis of LoRA within the Whisper encoder for SER. We employ a multifaceted interpretability toolkit that includes layer-wise contribution probing [12], Logit-Lens inspection [13], and representation analyses with SVD and CKA [14]. Our findings reveal two key mechanisms underlying LoRA’s success: a depth-specific delayed specialization strategy and a forward alignment, backward differentiation dynamic between LoRA’s low-rank matrices. These insights clarify how LoRA reorganizes encoder hierarchies and provide a foundation for designing more efficient and interpretable fine-tuning strategies for deep speech models.
@@ 2.1 Mechanistic Interpretability
Mechanistic interpretability (MI) is an emerging field that seeks to reverse-engineer the internal computations of deep networks to understand the algorithms they have learned [15]. In the domain of natural language processing (NLP), MI studies have shed light on how large language models (LLMs) encode complex behaviors like arithmetic and reasoning [16, 17]. For instance, Csordás et al. investigated the efficient use of network depth, revealing that in deep LLMs, the latter half of the network is often underutilized, primarily refining the output probability distribution rather than performing novel computations [12]. Building on this analytical framework, the ‘Backward Lens’ study extended it to the backward pass, revealing a two-phase ’imprint and shift’ mechanism for how models store new knowledge in their feed-forward layers [18]. Recent multimodal efforts, such as AudioLens [19], have further applied these principles to understand auditory perception in large audio-language models. However, within the speech processing community, MI remains limited, with most work focusing on probing intermediate representations rather than the intricate adaptation mechanisms of fine-tuning strategies like LoRA.
@@ 2.2 Speech Emotion Recognition
SER has evolved significantly, progressing from systems based on hand-crafted features [20] to those leveraging powerful self-supervised encoders like Wav2Vec2 [2] and HuBERT [3], and more recently, large pre-trained models such as Whisper [1]. This evolution has led to notable improvements in cross-speaker and cross-domain robustness [9]. Task-specific fine-tuning approaches, including those that incorporate metadata-enhanced strategies [21] and domain generalization for Whisper [10], have shown promising results. Nevertheless, the substantial computational and storage costs associated with full fine-tuning necessitate the exploration of more resource-efficient alternatives [22].
@@ 2.3 Parameter-Efficient Fine-Tuning
Parameter-Efficient Fine-Tuning methods [4, 5] offer a compelling solution by adapting frozen backbones with a minimal number of trainable parameters. Popular approaches include adapters, prefix-tuning, and prompt-tuning, among which LoRA [6] has demonstrated exceptional effectiveness. While adapters may introduce inference latency and prefix-tuning can be unstable for speech tasks, LoRA achieves remarkable efficiency without sacrificing performance by injecting low-rank updates. Its efficacy has been confirmed in various applications, including Whisper for ASR [8], yet the precise inner adaptation dynamics of LoRA for complex tasks like SER remain largely unexplored.
@@ 2.4 Research Gap
In summary, existing literature has made significant strides in three distinct areas: mechanistic interpretability in NLP, the application of large pretrained encoders for SER, and the development of PEFT methods for speech. However, the intersection of these fields, specifically the mechanistic interpretability of LoRA’s adaptation in speech emotion recognition, has not been systematically investigated. This work is dedicated to bridging this critical research gap by analyzing how LoRA reshapes the representational hierarchies and optimization dynamics within the Whisper encoder for the SER task.
3 Experimental Setup
SER is formulated as a four-class task (anger, happiness, neutrality, sadness) with WhisperForAudioClassification as the backbone. We conduct experiments on the IEMOCAP [23] dataset using a speaker-independent 10-fold cross-validation. Unweighted recall (UAR) and weighted recall (WAR) serve as our evaluation metrics; UAR provides a robust measure for imbalanced datasets, while WAR reflects overall accuracy. For our mechanistic analysis, we use the Whisper-large-v2 encoder and the NNsight [24] library, with analysis based on a hierarchical sample of 100 examples from a validation set (25 from each emotion class).
@@ 3.1 LoRA Fine-tuning
We adopt LoRA as our primary fine-tuning method, which approximates weight updates via a low-rank decomposition, $\Delta W=BA$. Only the low-rank matrices $A\in\mathbb{R}^{r\times d}$ and $B\in\mathbb{R}^{d\times r}$ are trainable, significantly reducing the number of parameters. We attach LoRA modules to the attention projections, a standard and effective strategy. Unless otherwise noted, we use a configuration of r=32 and apply a dropout of 0.1, with a trainable classification head.
@@ 3.3 Logit-Lens Inspection
As shown in Figure 2a, the two models exhibit distinct depth-wise alignment patterns. The frozen encoder shows volatile KL curves in early layers, drops to a minimum near layer 27, then rebounds at the top. This suggests that while Whisper possesses latent emotion-related capacity, achieving strongest emotional expression at the KL minimum, the original ASR objective creates inherent instability requiring downstream classifier compensation.
@@
3.4 Rank-Based Representation Analysis
Our t-SNE analysis across different LoRA ranks ($r=8,32,64$) reveals a systematic improvement in emotion clustering quality with increasing rank capacity. As shown in Figure  2b, deeper layers with $r=64$ achieve the most distinct and well-separated emotion boundaries, while maintaining a consistent depth-wise evolution from mixed to specialized representations. This pattern reinforces our delayed decision-making hypothesis: while the fundamental specialization mechanism remains consistent across ranks, LoRA’s representational capacity directly determines the final clustering quality and emotion separability. Low-rank configurations ($r=8$) demonstrate a fundamental limitation, exhibiting persistent inter-class overlap even in the deepest layers, suggesting minimal rank capacity constrains the model’s ability to achieve complete emotion separation.
@@ 3.5 Singular Value Decomposition Analysis
Fig. 3: Analysis of LoRA’s internal dynamics. (a)-(d): SVD analysis. (a, c) are for LoRAA and LoRAB activations, while (b, d) are for their gradients. (e), (f): CKA similarity for activations and gradients.
To investigate the intrinsic dimensionality and learning dynamics of LoRA, we conduct a comprehensive SVD analysis on its low-rank matrices across all layers. The “LoRA (mean)” is computed by averaging the singular values from all attention heads and all layers, providing a global perspective on its learning dynamics. By comparing a trained LoRA model against a randomly initialized counterpart, we analyze how each component concentrates energy within its dominant singular values. This analysis reveals a sophisticated division of labor between LoRA’s matrices, which develop specialized, complementary roles for compression and reconstruction.
@@
[…]
This complementary learning architecture—where LoRAA acts as an information compressor and LoRAB as a precision reconstructor—is a key to LoRA’s exceptional efficiency. This synergistic specialization ensures both aggressive forward compression and preserved gradient dimensionality, validating that LoRA operates within the intrinsic low-dimensional manifold of the adaptation task.
@@ 3.6 Centered Kernel Alignment Analysis
To quantify the representational similarity between LoRA components, we employ CKA. Unlike simple metrics such as cosine similarity, CKA provides a robust, normalized score invariant to linear transformations, making it ideal for comparing neural network representations. High values suggest the components capture similar representational structures, while low or negative values indicate dissimilar or inversely related structures
@@
where $\mathbf{K}$ and $\mathbf{L}$ are kernel matrices derived from representations $\mathbf{X}$ and $\mathbf{Y}$, and HSIC is the Hilbert-Schmidt Independence Criterion.
As depicted in Figure 3ef, our analysis reveals a crucial asymmetry in LoRA’s learning dynamics. The forward passes of the LoRA components consistently maintain high representational similarity (0.8-1.0), indicating that LoRA’s forward pass acts as a coherent information flow. In stark contrast, their corresponding gradient similarities exhibit a substantial variability, ranging from high positive to negative correlations. The presence of negative correlations is particularly revealing, as it indicates an antagonistic optimization dynamic, where LoRAA and LoRAB actively pull in opposite directions in certain layers during the backward pass.
This finding elucidates a fundamental forward alignment, backward differentiation mechanism. It shows that LoRA achieves representational consistency during inference (forward) while maintaining specialized, fine-grained optimization roles during training (backward). Our analysis suggests that LoRAA functions primarily as a feature compressor, guided by these differentiated backward signals, while LoRAB acts as a precision reconstructor. This synergistic specialization enables LoRA to achieve exceptional parameter efficiency without sacrificing the model’s overall expressivity.
@@ 4 CONCLUSION
This paper presents the first systematic mechanistic analysis of LoRA in the Whisper encoder for SER. Our analysis reveals two key mechanisms for its superior performance: “delayed specialization”, which preserves stable representations in early layers for decisive late-stage consolidation, and a “forward alignment, backward differentiation” dynamic between LoRA’s matrices. These insights not only clarify LoRA’s effectiveness but also establish a theoretical foundation for understanding low-rank fine-tuning in deep speech models. By demonstrating how a small number of parameters can fundamentally restructure the information flow, our work paves the way for the design of more efficient and interpretable adaptation strategies. Future work will investigate whether these dynamics generalize to other speech tasks, such as multilingual emotion recognition or speaker identification, and whether they are consistent across different large-scale pre-trained models. Code is available at https://github.com/harryporry77/Behind-the-Scenes

### v3

@@ Abstract
Large pre-trained speech models such as Whisper offer strong generalization but pose significant challenges for resource-efficient adaptation. Low-Rank Adaptation (LoRA) has become a popular parameter-efficient fine-tuning method, yet its underlying mechanisms in speech tasks remain poorly understood. In this work, we conduct the first systematic mechanistic interpretability study of LoRA within the Whisper encoder for speech emotion recognition (SER). Using a suite of analytical tools, including layer contribution probing, logit-lens inspection, and representational similarity via singular value decomposition (SVD) and centered kernel alignment (CKA), we reveal two key mechanisms: a delayed specialization process that preserves general features in early layers before consolidating task-specific information, and a forward alignment, backward differentiation dynamic between LoRA’s matrices. Our findings clarify how LoRA reshapes encoder hierarchies, providing both empirical insights and a deeper mechanistic understanding to design efficient and interpretable adaptation strategies in large speech models.
@@ 1 Introduction
Large-scale pre-trained models have fundamentally transformed speech and language processing. Models such as Whisper [1], Wav2Vec2 [2], and HuBERT [3] provide general-purpose representations that are widely adapted to diverse tasks, from automatic speech recognition (ASR) to SER. Yet this versatility comes at a cost: the heavy computational and storage demands of full fine-tuning often hinder their deployment in resource-constrained settings. This tension has motivated the development of Parameter-Efficient Fine-Tuning (PEFT) methods [4, 5], which aim to adapt large encoders with minimal additional parameters.
Among PEFT approaches, LoRA [6] has demonstrated strong effectiveness across multiple modalities. By injecting trainable low-rank matrices into frozen backbones, LoRA drastically reduces training costs while preserving the representational power of pre-trained models. Despite this practical success, most previous studies have focused on reporting performance improvements, while the underlying mechanisms behind LoRA’s effectiveness remain underexplored. Addressing this interpretability gap is critical for moving beyond empirical gains toward principled design of efficient fine-tuning strategies.
SER is not only a practically important task in affective computing but also a compelling domain for investigating parameter-efficient adaptation. Unlike ASR, which relies primarily on lexical information, SER is a classification task that depends on prosodic and paralinguistic cues reflecting high-level affective states [7], which makes it well suited to analyze how LoRA’s low-rank adaptation reshapes discriminative representations. The successful adaptation of Whisper [8] and Wav2Vec2 [9] to SER further underscores the dual importance of enhancing performance and understanding underlying mechanisms in affective speech modeling.
Despite the growing popularity of parameter-efficient tuning, the internal mechanisms by which LoRA reshapes large speech encoders remain poorly understood, particularly for classification-oriented tasks such as SER. To bridge this gap, we present a systematic mechanistic study of LoRA within the Whisper encoder, leveraging probing and representational analyses to examine its adaptation process. Our code is publicly available [^1] https://github.com/harryporry77/Behind-the-Scenes.
@@ 2.1 Mechanistic Interpretability
Mechanistic interpretability (MI) is an emerging field that seeks to reverse-engineer the internal computations of deep networks to understand the algorithms they have learned [10]. In the domain of natural language processing (NLP), MI studies have shed light on how large language models (LLMs) encode complex behaviors such as arithmetic and reasoning [11, 12]. For instance, Csordás et al. investigated the efficient use of network depth, revealing that in deep LLMs, the latter half of the network is often underutilized, primarily refining the output probability distribution rather than performing novel computations [13]. Building on this analytical framework, the ‘Backward Lens’ study extended it to the backward pass, revealing a two-phase ’imprint and shift’ mechanism for how models store new knowledge in their feed-forward layers [14]. Recent multimodal efforts, such as AudioLens [15], have further applied these principles to understand auditory perception in large audio-language models. However, within the speech processing community, MI remains limited, with most work focusing on probing intermediate representations rather than the intricate adaptation mechanisms of fine-tuning strategies like LoRA.
@@ 2.2 Speech Emotion Recognition
SER has evolved significantly, progressing from systems based on hand-crafted features [16] to those leveraging powerful self-supervised encoders like Wav2Vec2 [2] and HuBERT [3], and more recently, large pre-trained models such as Whisper [1]. This evolution has led to notable improvements in cross-speaker and cross-domain robustness [7, 17]. Task-specific fine-tuning approaches, including those that incorporate metadata-enhanced strategies [18] and domain generalization for Whisper [8], have shown promising results. Nevertheless, the substantial computational and storage costs associated with full fine-tuning necessitate the exploration of more resource-efficient alternatives [19, 20].
@@ 2.3 Parameter-Efficient Fine-Tuning
PEFT methods [4, 5] offer a compelling solution by adapting frozen backbones with a minimal number of trainable parameters. Popular approaches include adapters, prefix-tuning, and prompt-tuning, among which LoRA [6] has demonstrated exceptional effectiveness. While adapters may introduce inference latency and prefix-tuning can be unstable for speech tasks, LoRA achieves remarkable efficiency without sacrificing performance by injecting low-rank updates. Its efficacy has been confirmed in various applications, including Whisper for ASR [21], yet the precise inner adaptation dynamics of LoRA for complex tasks like SER remain largely unexplored.
@@ 2.4 Research Gap
In summary, existing literature has made significant strides in three distinct areas: mechanistic interpretability in NLP, the application of large pretrained encoders for SER, and the development of PEFT methods for speech. However, the intersection of these fields, specifically the mechanistic interpretability of LoRA’s adaptation in speech emotion recognition, has not been systematically investigated. This work is dedicated to bridging this critical research gap by analyzing how LoRA reshapes the representational hierarchies and optimization dynamics within the Whisper for the SER.
3 Experiment
SER is formulated as a 4-class task (anger, happiness, neutrality, sadness) with WhisperForAudioClassification as the backbone. We conduct experiments on the IEMOCAP [22] dataset using a standard speaker-independent 10-fold cross-validation. Unweighted recall (UAR) and weighted recall (WAR) serve as our evaluation metrics; UAR provides a robust measure for imbalanced datasets, while WAR reflects overall accuracy. For our mechanistic analysis, we use the Whisper-large-v2 encoder and the NNsight [23] library, with analysis based on a hierarchical sample of 100 examples from a validation set (25 from each emotion class). We freeze the entire Whisper encoder (including LayerNorm and positional embeddings). For LoRA, only $\{A,B\}$ and the classification head are trainable; for the Frozen-Encoder baseline, only the same head is trained.
@@ 3.1 LoRA Fine-tuning
[…]
@@ 3.3 Logit-Lens Inspection
As shown in Fig. 2a, the two models exhibit distinct depth-wise alignment patterns. The frozen encoder shows volatile KL curves in early layers, drops to a minimum near layer 27, then rebounds at the top. This suggests that while Whisper possesses latent emotion-related capacity, achieving strongest emotional expression at the KL minimum, the original ASR objective creates inherent instability requiring downstream classifier compensation.
@@
3.4 Rank-Based Representation
Our t-SNE analysis across different LoRA ranks ($r=8,32,64$) reveals a systematic improvement in emotion clustering quality with increasing rank capacity. As shown in Fig.  2b, deeper layers with $r=64$ achieve the most distinct and well-separated emotion boundaries, while maintaining a consistent depth-wise evolution from mixed to specialized representations. This pattern reinforces our delayed decision-making hypothesis: while the fundamental specialization mechanism remains consistent across ranks, LoRA’s representational capacity directly determines the final clustering quality and emotion separability. Low-rank configurations ($r=8$) demonstrate a fundamental limitation, exhibiting persistent inter-class overlap even in the deepest layers, suggesting minimal rank capacity constrains the model’s ability to achieve complete emotion separation.
@@ 3.5 Singular Value Decomposition Analysis
Fig. 3: Analysis of LoRA’s internal dynamics. (a–d) SVD-based effective-rank curves for LoRAA/LoRAB activations and their gradients, comparing trained LoRA (solid) with random initialization (dashed). The y-axis is cumulative energy ratio and the x-axis is the number of singular components. (e) CKA between LoRAA and LoRAB for activations and gradients.
To investigate the intrinsic dimensionality and learning dynamics of LoRA, we perform SVD on the activation matrices produced by LoRAA and LoRAB, and on their corresponding gradient matrices, across all layers. The “LoRA (mean)” is computed by averaging the spectra across layers (and attention heads when applicable), providing a global perspective on learning dynamics. Throughout, we use an energy-based effective-rank view: we report how many singular components are required to reach a fixed cumulative energy threshold (90%), and the reported values can be non-integer due to averaging across layers. By comparing a trained LoRA model against a randomly initialized counterpart, we analyze how each component concentrates energy within its dominant singular components. This analysis suggests complementary roles between LoRA’s components for compression and reconstruction.
@@
This complementary learning behavior, where LoRAA concentrates energy into a low-dimensional subspace while LoRAB preserves richer gradient diversity, helps explain LoRA’s parameter efficiency. Overall, the results are consistent with the adaptation being effectively low-dimensional under this task setting.
@@ 3.6 Centered Kernel Alignment Analysis
To quantify representational similarity between LoRA components, we employ CKA [25], a normalized alignment measure for comparing neural representations across layers. It is robust to isotropic scaling and orthogonal transformations of the representations, and is therefore more appropriate than raw cosine similarity for high-dimensional feature comparisons. We compute linear CKA layerwise for both forward activations and backward gradient signals.
@@
where $\mathbf{X}\in\mathbb{R}^{n\times d_{x}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{y}}$ are representation matrices with rows as examples and columns as features. $\mathbf{K}$ and $\mathbf{L}$ are centered Gram matrices derived from $\mathbf{X}$ and $\mathbf{Y}$, and HSIC [26] is a kernel-based dependence measure computed from these centered Gram matrices. For forward activations, $\mathbf{X}$ and $\mathbf{Y}$ correspond to LoRAA/LoRAB output activations at the same layer. For backward signals, $\mathbf{X}$ and $\mathbf{Y}$ correspond to the gradients of the loss with respect to these LoRA output activations.
As depicted in Fig. 3e, LoRA exhibits a clear forward and backward asymmetry. Forward activations of LoRAA and LoRAB remain highly aligned (0.8–1.0) across layers, indicating that the two components induce very similar representational structure during the forward pass. In contrast, gradient alignment is strongly layer-dependent and varies substantially, suggesting that the two components can receive differentiated optimization signals even when their forward representations are similar.
SVD and CKA are complementary: the former characterizes energy concentration within each component, and the latter quantifies layerwise alignment between components. Together they indicate compression by LoRAA and reconstruction by LoRAB with forward alignment and backward differentiated dynamics.
@@ 4 CONCLUSION
This paper provides the first systematic mechanistic analysis of LoRA’s role within the Whisper encoder for SER. Our findings reveal two key mechanisms that contribute to LoRA’s superior performance: a delayed specialization strategy, which maintains stable representations in early layers to enable decisive late-stage consolidation, and a forward alignment, backward differentiation dynamic between LoRA’s matrices. These insights not only elucidate LoRA’s effectiveness but also offer a deeper understanding of how low-rank fine-tuning methods fundamentally restructure the information flow within deep speech models. Our work establishes a foundation for designing more efficient and interpretable adaptation strategies across various speech tasks, and we believe our analysis framework can serve as a blueprint for investigating whether these dynamics generalize to other models and datasets.